OpenStudy (anonymous):

Solve for x: 2/5 + 3/5x = x+5/10

OpenStudy (anonymous):

x=-1/4

OpenStudy (anonymous):

4+6x=10x+5 -4x=1 x=-1/4

OpenStudy (anonymous):

How do you get 10x +5 on the right?

OpenStudy (anonymous):

oh sorry. Multiply both sides by 10 so you can get rid of the fractions :)

OpenStudy (anonymous):

multiply the whole equation by 10, so: 4 + 6x = 10x + 5 so: 4x = -1 so: x = -1/4

OpenStudy (anonymous):

okay thank you, it was simple